OpenStudy (anonymous):

If f(x) = −4x − 2, find f(−3) 10 −14 12 −10

OpenStudy (anonymous):

f(x) = -4x -2 f(-3) = -4(-3) -2 = .................. see we have just simply substituted x=-3 into the equation given?

OpenStudy (anonymous):

yes(:

OpenStudy (anonymous):

ok, now tell me what the answer is if you have figured it out?

OpenStudy (anonymous):

-228

OpenStudy (anonymous):

ok this is our equation f(x) = -4x -2 it's the same as saying y = -4x -2 and we want to substitute x=-3 into the equation, so we simply replace the x in the above equation by -3 y=-4(-3) -2 y= ? (you can do this)

i dont know what to do ):